Can some one help me with this?
One-third of a number is 5 less than half of the same number. What is the number?

Answer by mukhopadhyay(490)

Let the number be x
x/3 = x/2 - 5
=>x/2 - x/3 = 5
=>x/6 = 5
=>x = 30
The answer is 30